OpenStudy (owlstudies):

PLEASE HELP: An employee at a major corporation has $2,000 to invest in her 401K. She wants to invest part of it in a mutual fund, part in company stock and part in bonds. She has decided that she wants to invest twice as much in the mutual fund as she invests in company stock. the first three months of investment, the mutual fund increases in value by 4.5%, the bonds increase in value by 2.6% and the company stock falls in value by 0.2%. As a result of these changes, the value in her investment increased by a total of $58. Based on this information, how much money did she initially invest in the mutual fund, bonds, and company stock?

OpenStudy (retireed):

Something is missing in this sentence... "As a result these the value her investment initially increased by a total of on this information, " her investment initially increased by a total of ______ (what, there needs to be a value here) The sentence at the point "of on" makes no sense. As a result these the value her investment initially increased by a total of on this information, how much money did she invest in the mutual fund, bonds and company stock?

OpenStudy (owlstudies):

I'm sorry i don't know what happened. It's "as a result of these changes, the value in her investment increased by a total of $58. Based on this information, how much money did she initially invest in the mutual fund, bonds, and company stock?"

OpenStudy (retireed):

Thanks for the correction... now the problem can be solved, but it will take me a while to solve it. I always like to try and solve it before I try to help some one, so I know how to do it myself. Since you are here, we can try to solve it together...

OpenStudy (retireed):

First we need to define the variables... Invested $2000 MF = mutal funds B = bonds and S =stocks MF = 2(S) B = 2000 - MF - S since MF = 2S B = 2000 - 3S now it starts to get difficult....

OpenStudy (retireed):

the mutual fund increases in value by 4.5%, the bonds increase in value by 2.6% and the company stock falls in value by 0.2%. As a result of these changes, the value in her investment increased by a total of $58. Based on this information, how much money did she initially invest in the mutual fund, bonds, and company stock? 2058 = MF(0.045) + B(0.026) + S(-0.002) I think this looks correct We all ready have B = 2000 - 3S and MF = 2S You can do the rest. Plug the last two equations into the one above them 2058 = 2S(0.045) + (2000-3S)(0.026) + S(-0.002) Now solve this for S, then double it to get the MF finally solve for B. I will leave the calculation to you...

OpenStudy (owlstudies):

ohhh ok i get it now. thank you so much!

OpenStudy (retireed):

great!